OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in standard form

OpenStudy (anonymous):

3=-6+b 9=b y=x+9

OpenStudy (anonymous):

Sorry I messed up the first time.

OpenStudy (anonymous):

Thanks!